# Interaction-Augmented Instruction: Modeling the Synergy of Prompts and Interactions in Human-GenAI Collaboration

### v1

@@ Abstract.
Text prompt is the most common way for human-generative AI (GenAI) communication. Though convenient, it is challenging to convey fine-grained and referential intent. One promising solution is to combine text prompts with precise GUI interactions, like brushing and clicking. However, there lacks a formal model to capture synergistic designs between prompts and interactions, hindering their comparison and innovation. To fill this gap, via an iterative and deductive process, we develop the Interaction-Augmented Instruction (IAI) model, a compact entity–relation graph formalizing how the combination of interactions and text prompts enhances human-GenAI communication. With the model, we distill twelve recurring and composable atomic interaction paradigms from prior tools, verifying our model’s capability to facilitate systematic design characterization and comparison. Four usage scenarios further demonstrate the model’s utility in applying, refining, and extending these paradigms. These results illustrate the IAI model’s descriptive, discriminative, and generative power for shaping future GenAI systems.
Generative AI, Text Prompt, Interaction, Human-GenAI Collaboration
copyright: rightsretainedconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 5837ccs: Human-centered computing Interaction paradigms
@@ 1. Introduction
Generative AI (GenAI) has rapidly become a general-purpose technology, enabling new intelligent applications across domains such as design (Shen et al., 2025a), education (Law, 2024), and data analysis (Inala et al., 2024). Modern large language and multimodal models can interpret open-ended text prompts and generate diverse outputs (e.g., text, image, and data analysis dashboards), making the text prompt interface the dominant paradigm for human–AI communication (Shen et al., 2023). By allowing users to express complex intents in natural language, text prompts offer great flexibility, allowing users “ask anything” and receive tailored responses.
@@
Despite the convenience, free-form text prompts are often insufficient to convey fine-grained and nuanced human intent when precise instructions are required. This is due to the ambiguity and the coarse nature of natural language (Shen et al., 2025b; Riche et al., 2025; Subramonyam et al., 2024). For instance, in image editing, a user might want to modify “the upper-right flower petal”, but a text prompt alone is hard to accurately localize that target visual element. Likewise, in coding, asking “write a loop to process this data” could imply different iteration strategies. Correspondingly, in practice, users often struggle to articulate sufficient detail or to break down complex goals into a single prompt, resulting in suboptimal or hard-to-control AI behavior.
To address these limitations, a promising strategy is to combine general-purpose but imprecise text prompts with dedicated interactions (e.g., clicks, drags, brushes, etc.) through graphical user interfaces (GUIs), as shown in Fig. 1-A. The synergy enables both flexible and fine-grained control by users. For example, OpenAI Canvas (OpenAI, 2024) and Cursor (Anysphere, 2023) let users select a specific text segment or code block and then issue a focused prompt that applies only to that context. Tableau AI (Tableau, 2024) uses LLMs to suggest context-aware follow-up questions on GUI for user selection during data exploration. Research prototypes following this strategy have also emerged, exploiting the strengths of both natural language and direct-manipulation input for diverse tasks, such as coding (Feng et al., 2024), writing (Laban et al., 2024), data visualization (Xie et al., 2025), and image editing (Liu et al., 2025). Despite the wide and diverse applications, the strategy has not been formalized as a comprehensive interaction model to guide related research and practices. As a pioneering study in the HCI field points out (Beaudouin-Lafon, 2004), these “point-like” research demonstrations of the strategy are often insufficient due to the challenges of transforming task-specific demonstrations to broader real-world applications and understanding the advancement of related techniques. It is essential to propose general interaction models to facilitate the careful examination of existing designs (descriptive power), the comparison between design variations (discriminative power), and the consequent innovation of new interaction designs for emerging tasks (generative power(Beaudouin-Lafon, 2004, 2000; Liao and Xiao, 2023).
@@
To fill this gap, through a deductive and iterative process, we propose the Interaction-Augmented Instruction (IAI) model: a compact entity–relation graph that makes explicit how precise GUI interactions and free-form prompts are composed into the executable instructions consumed by GenAI (Fig. 1-B). The model comprises six entities: Human (H), Interaction (I), Text Prompt (T), Augmented Instruction (Aug), GenAI (G), and Artifact (A). By consolidating candidate roles into this minimal set, the model highlights Aug as the single explicit input to GenAI, thereby simplifying comparison across paradigms and clarifying how different interfaces combine T, I, and A to construct Aug. We further systematically enumerated all pairwise relations and preserved only those that are semantically meaningful, discriminative, and interpretable (Table 1).
Building on this foundation, to examine if our IAI model can be generally applied to describe and differentiate existing GenAI systems’ interfaces, we revisit a curated corpus of GenAI-powered interactive systems and map each system’s interaction workflow to one or more directed paradigm graphs (Fig. 1-C). From the corpus, we distill twelve frequent and composable atomic paradigms organized by interaction timing (pre- or post- GenAI invocation) and resource availability (prompt-only vs. artifact-grounded). These paradigm graphs based on the IAI model serve as high-level abstractions of interaction workflows, enabling tools to be clustered and compared.
The IAI model provides not only a theoretical lens but also a practical framework for future interface design. The twelve design paradigms derived from the IAI model can be applied, refined, and extended for creating new interfaces that meet diverse scenario demands. Unexplored interaction design paradigms are also revealed from comparing the existing paradigms with the entire IAI model. To demonstrate the power of generating new interface design and novel interaction paradigms, we present four usage scenarios that bridge conceptual models with actionable design processes (Fig. 1-D). These usage scenarios show how the IAI model can guide the iterative enhancement of existing tools by chaining or editing existing design paradigms, while also supporting the adaptation of established paradigms and the derivation of new atomic paradigms for emerging scenarios. Looking forward, we envision the IAI model as a foundation for advancing human–GenAI communication, opening new opportunities to design more transparent, controllable, and creative GenAI systems. To summarize, our work makes three main contributions:
@@ 2.1. Human–GenAI Communication
Free-form text prompts remain the dominant way for human–GenAI communication due to their flexible, expressive, and low-friction (Wu et al., 2022; Chung and Adar, 2023). However, prompt-only interaction has well-documented limitations: natural language is often ambiguous, underspecified, or ill-structured for task-specific operations, which increases iteration cost and reduces control (Riche et al., 2025; Subramonyam et al., 2024; Shen et al., 2025b). Empirical studies highlight recurring challenges such as prompt formulation, disambiguation, and intent steering—for example, users struggle to localize visual elements precisely through language alone (Masson et al., 2024; Fok et al., 2024) or to specify the exact semantic behavior desired in code edits (Feng et al., 2024; Yen et al., 2024).
@@
Another line of interface design has been to encapsulate GenAI capabilities within backend agents or domain-specific tools (Bansal et al., 2024). In such designs, the system exposes a constrained interface (e.g., widgets, templates, or menus) (Lin et al., 2023; Masson et al., 2025; Wang et al., 2025) or totally automate the process while GenAI executes domain logic behind the scenes. For example, Textoshop (Masson et al., 2025) allows text editing entirely through drawing software-like interactions, with prompts handled internally; fully automated agents similarly hide prompting from the user (Shen et al., 2024a; Zhao et al., 2024). This lowers user burden but constrains openness and the expressive flexibility of prompting.
@@
A more flexible and increasingly prevalent approach is to combine prompting with GUI interactions—what we call interaction-augmented instructions. Here, prompts are enriched by targeted user actions such as clicking, brushing, sketching, or selecting (Gao et al., 2024; Shen et al., 2025b; Hu et al., 2025). These interactions may occur before the GenAI invocation, e.g., selecting elements to constrain the scope of a prompt (Xie et al., 2024; Masson et al., 2024; Wang et al., 2024c; Chen et al., 2024; Fok et al., 2024; Liu et al., 2025, 2023a), organizing multiple prompts into structured forms such as trees or graphs (Feng et al., 2024; Yen et al., 2024; Wu et al., 2022; Gmeiner et al., 2025), sketching to guide content generation (Lin et al., 2025; Kim et al., 2023a; Nguyen et al., 2025; Zhou et al., 2025), annotating text to extend context or highlight specific information (Shen et al., 2024b), or after it, e.g., clarifying user intent through follow-up queries (Bursztyn et al., 2021; Baek et al., 2023; Andrews et al., 2024; Ma et al., 2024; Gmeiner et al., 2025), presenting suggestions for human confirmation (Laban et al., 2024; Wu et al., 2023; Angert et al., 2023), converting prompt into structured GUI components (Zhang et al., 2025; Cai et al., 2024), or enabling direct post-generation artifact manipulation (Xie et al., 2025; Vaithilingam et al., 2024; Suh et al., 2024; Gero et al., 2022; Petridis et al., 2024; Jiang et al., 2023). Together, these paradigms illustrate diverse strategies for balancing the openness of prompting with the precision of interaction.
@@ 2.2. Human–GenAI Interaction Models and Design Patterns
Interaction design in modern UIs has evolved from early WIMP and direct-manipulation paradigms to richer post-WIMP models that foreground instruments and objects over low-level commands (Van Dam, 1997; Lee et al., 2021). Direct-manipulation interfaces emphasized immediate, visible control—users act on representations of domain objects and observe one-to-one effects (Shneiderman, 1983). Building on this view, Beaudouin-Lafon’s Instrumental Interaction framed instruments as mediators between user actions and artifacts, where each instrumented action maps directly to a target operation (Beaudouin-Lafon, 2000; Beaudouin-Lafon and Mackay, 2000). Subsequent work introduced “substrates” as “places for interaction”, a mechanism to combine power and simplicity by structuring where and how instruments apply in complex interfaces (Mackay and Beaudouin-Lafon, 2025).
These classical models are well suited to deterministic manipulation, but they assume a one-to-one relation between an interaction and its effect on an artifact. Generative systems violate that assumption: a single high-level instruction (e.g., a free-form text prompt) can trigger multiple, nonlocal transformations across an artifact or even across multiple artifacts. This multiplicity and indirection require rethinking how instruments, intentions, and objects are modeled in an era of GenAI.
Recent HCI work has begun to explore this space. A series of surveys and taxonomies catalog interaction modalities and recurring patterns. For example, Lehmann et al.  (Lehmann and Buschek, 2024) analyze how UI affordances mediate access to model capabilities, Luera et al.  (Luera et al., 2024) review input modalities across GenAI applications, Gao et al.  (Gao et al., 2024) propose a taxonomy of human–LLM communication modes, and Hu et al.  (Hu et al., 2025) identify general paradigms, i.e., GenAI as medium, tool, partner, and mediator. Shen et al.  (Shen et al., 2025b) focus specifically on interaction-augmented instructions, outlining four core purposes for interaction, i.e., restricting, expanding, organizing, and refining prompts. Complementing these descriptive efforts, Riche et al.  (Riche et al., 2025) introduce AI-Instruments, extending instrumental interaction (Beaudouin-Lafon, 2000) to generative settings by making prompts themselves manipulable objects and articulating principles of reification, reflection, and grounding.
Prior work has laid important groundwork but remains limited in different ways. Taxonomies of GenAI interaction paradigms (Gao et al., 2024; Hu et al., 2025) primarily offer descriptive power, enumerating recurring patterns without providing a mechanism to compare or compose them. In contrast, AI-Instruments (Riche et al., 2025), grounded in instrumental interaction (Beaudouin-Lafon, 2000), articulate generative principles for reifying and reflecting user intent, but they stop short of offering descriptive and discriminative power across diverse system interfaces. This paper is intended to fill that gap by providing a formal model for describing, discriminating, and generating interaction-augmented GenAI paradigms.

### v2

@@ Abstract.
Text prompt is the most common way for human-generative AI (GenAI) communication. Though convenient, it is challenging to convey fine-grained and referential intent. One promising solution is to combine text prompts with precise GUI interactions, like brushing and clicking. However, there lacks a formal model to capture synergistic designs between prompts and interactions, hindering their comparison and innovation. To fill this gap, via an iterative and deductive process, we develop the Interaction-Augmented Instruction (IAI) model, a compact entity–relation graph formalizing how the combination of interactions and text prompts enhances human-GenAI communication. With the model, we distill twelve recurring and composable atomic interaction paradigms from prior tools, verifying our model’s capability to facilitate systematic design characterization and comparison. Four usage scenarios further demonstrate the model’s utility in applying, refining, and innovating these paradigms. These results illustrate the IAI model’s descriptive, discriminative, and generative power for shaping future GenAI systems.
Generative AI, Prompt, Interaction, Human-GenAI Collaboration
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790505isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interaction paradigms
@@ 1. Introduction
Generative AI (GenAI) has rapidly become a general-purpose technology, enabling new intelligent applications across domains such as design (Shen et al., 2025b), education (Law, 2024), and data analysis (Inala et al., 2024). Modern large language and multimodal models can interpret open-ended text prompts and generate diverse outputs (e.g., text, image, video), making the text prompt interface the dominant paradigm for human–AI communication (Shen et al., 2023). Yet free-form text is often too ambiguous and coarse to express fine-grained or nuanced intent (Shen et al., 2025a; Riche et al., 2025; Subramonyam et al., 2024). For instance, in image editing, a user might want to modify “the upper-right flower petal”, but a text prompt alone is hard to accurately localize that target visual element. Likewise, in coding, asking “write a loop to process this data” could imply different iteration strategies. Correspondingly, in practice, users often struggle to articulate sufficient detail or to break down complex goals into a single prompt.
To address these limitations, a promising strategy is to combine general-purpose but imprecise text prompts with dedicated interactions (e.g., clicks, drags, brushes, etc.) through graphical user interfaces (GUIs), as shown in Fig. 1-A. The synergy enables both flexible and fine-grained user control.   For example, consider Jason, a creative director designing a stylized magazine cover portrait. His intent is clear: “I need a stylized portrait for a magazine cover, with high fashion and artistic lighting”. Compared with the basic prompt-only workflow (Fig. 2-a), this task can be realized through a spectrum of interface designs (Fig. 2 b–f) that combine prompts and interactions in distinct ways. Concretely, some interfaces let him refine or expand the text prompt before generation (b), for instance, selecting “artistic lighting” for fine-tuning or switching to a “minimal black-and-white” style via UI controls. This functionality speeds up exploration but still assumes all intent can be expressed in language (Brade et al., 2023; Almeda et al., 2024). To go beyond language, Jason might manually sketch a portrait outline or brush key regions, merging these non-linguistic signals with the prompt for more precise spatial and referential control (c) (Lin et al., 2025; Kim et al., 2023b). When an initial render already exists, he may directly draw a necklace on the portrait, with the system incorporating both the prompt and the manipulated artifact content to enable precise, auditable edits grounded in the actual image (d) (Masson et al., 2024; Liu et al., 2025). After generating an initial image, some systems propose interactive controls (e), such as sliders for keyword (e.g., stylized” and “high fashion”) attention or color palettes for finer “artistic lighting”, allowing Jason to steer subsequent generations without rewriting text (Wang et al., 2024d; Gmeiner et al., 2025). Others analyze the generated artifact and propose structured options (f): Jason can choose to generate cover text, select the type, style, and length, and assemble a structured instruction for a magazine tagline (Zhang et al., 2023; Kim et al., 2023a).
Figure 2.   A unified image-generation task realized with distinct interaction-augmented designs: (a) prompt-only baseline; (b) prompt refinement via interaction; (c) combining prompt with non-linguistic interaction signals; (d) grounding instructions in artifact manipulation; (e) post-generation prompt control widgets; (f) artifact-anchored structured follow-ups.
Although these designs support the same creative goal, they differ fundamentally in various dimensions, such as where user intent is specified (e.g., text, interaction, artifact), when interactions occur (pre- or post-model invocation), and what information is ultimately delivered to GenAI. Those structural differences affect referential fidelity, controllability, provenance, and user experience. Furthermore, such strategies have not yet been formalized as a comprehensive interaction model to guide related research and practices. As a pioneering study in the HCI field notes (Beaudouin-Lafon, 2004), these “point-like” research demonstrations of the strategy are often insufficient due to the challenges of transforming task-specific demonstrations to broader real-world applications and understanding the advancement of related techniques. It is essential to propose general interaction models to facilitate the careful examination of existing designs (descriptive power), the comparison between design variations (discriminative power), and the consequent innovation of new interaction designs for emerging tasks (generative power(Beaudouin-Lafon, 2004, 2000; Liao and Xiao, 2023).
@@
To fill this gap, through a deductive and iterative process, we propose the Interaction-Augmented Instruction (IAI) model: a compact entity–relation graph that makes explicit how precise GUI interactions and free-form prompts are composed into the executable instructions consumed by GenAI (Fig. 1-B). The model comprises six entities: Human (H), Interaction (I), Text Prompt (T), Augmented Instruction (Aug), GenAI (G), and Artifact (A). By consolidating candidate roles into this minimal set, the model highlights Aug as the explicit input to GenAI, thereby simplifying comparison across paradigms and clarifying how different interfaces combine T, I, and A to construct Aug. We further systematically enumerated all pairwise relations and preserved only those that are semantically meaningful, discriminative, and interpretable (Tab. 1).
Building on this foundation, to examine if our IAI model can be generally applied to describe and differentiate existing GenAI systems’ interfaces, we revisit a curated corpus [^1] The annotation results: https://interaction-augmented-instruction.github.io/ of GenAI-powered interactive systems and map each system’s interaction workflow to one or more directed paradigm graphs (Fig. 1-C). From the corpus, we distill twelve frequent and composable atomic paradigms organized by interaction timing (pre- or post- GenAI invocation) and resource availability (prompt-only vs. artifact-grounded). These paradigm graphs based on the IAI model serve as high-level abstractions of interaction workflows, enabling tools to be clustered and compared.
The IAI model provides not only a theoretical lens but also a practical framework for future interface design. The twelve design paradigms derived from the IAI model can be applied, refined, and extended for creating new interfaces that meet diverse scenario demands. Unexplored interaction design paradigms are also revealed from comparing the existing paradigms with the entire IAI model. To demonstrate the power of generating new interface design and novel interaction paradigms, we present four usage scenarios that bridge conceptual models with actionable design processes (Fig. 1-D). These usage scenarios show how the IAI model can guide the iterative enhancement of existing tools by chaining or editing existing design paradigms, while also supporting the adaptation of established paradigms and the derivation of new atomic paradigms for emerging scenarios. Looking forward, we envision the IAI model as a foundation for advancing human–GenAI communication, opening new opportunities to design more transparent, controllable, and creative GenAI systems.
To summarize, our work makes three main contributions:
@@ 2.1. Human–GenAI Communication
Free-form text prompts remain the dominant way for human–GenAI communication due to their flexible, expressive, and low-friction (Wu et al., 2022; Chung and Adar, 2023). However, prompt-only interaction has well-documented limitations: natural language is often ambiguous, underspecified, or ill-structured for task-specific operations, which increases iteration cost and reduces control (Riche et al., 2025; Subramonyam et al., 2024; Shen et al., 2025a). Empirical studies highlight recurring challenges such as prompt formulation, disambiguation, and intent steering—for example, users struggle to localize visual elements precisely through language alone (Masson et al., 2024; Fok et al., 2024) or to specify the exact semantic behavior desired in code edits (Feng et al., 2024; Yen et al., 2024).
@@
Another line of interface design has been to encapsulate GenAI capabilities within backend agents or domain-specific tools (Bansal et al., 2024). In such designs, the system exposes a constrained interface (e.g., widgets, templates, or menus) (Lin et al., 2023; Masson et al., 2025; Wang et al., 2025, 2024b) or totally automate the process while GenAI executes domain logic behind the scenes (Shen et al., 2024c, b; Shao et al., 2025). For example, Textoshop (Masson et al., 2025) allows text editing entirely through drawing software-like interactions, with prompts handled internally; fully automated agents similarly hide prompting from the user (Shen et al., 2024b; Zhao et al., 2024). This lowers user burden but constrains openness and the expressive flexibility of prompting.
@@
A more flexible and increasingly prevalent approach is to combine prompting with GUI interactions—what we call interaction-augmented instructions. Here, prompts are enriched by targeted user actions such as clicking, brushing, sketching, or selecting (Gao et al., 2024; Shen et al., 2025a; Hu et al., 2025; Wang et al., 2023; Luo et al., 2025). These interactions may occur before the GenAI invocation, e.g., selecting elements to constrain the scope of a prompt (Xie et al., 2024; Masson et al., 2024; Wang et al., 2024a; Chen et al., 2024; Fok et al., 2024; Liu et al., 2025, 2023b; Ouyang et al., 2024; Zhang et al., 2025a), organizing multiple prompts into structured forms such as trees or graphs (Feng et al., 2024; Yen et al., 2024; Wu et al., 2022; Gmeiner et al., 2025), sketching to guide content generation (Lin et al., 2025; Kim et al., 2023b; Nguyen et al., 2025; Zhou et al., 2025), annotating text to extend context or highlight specific information (Shen et al., 2024a), or after it, e.g., clarifying user intent through follow-up queries (Bursztyn et al., 2021; Baek et al., 2023; Andrews et al., 2024; Ma et al., 2024; Gmeiner et al., 2025), presenting suggestions for human confirmation (Laban et al., 2024; Wu et al., 2023; Angert et al., 2023; Gao et al., 2025), converting prompt into structured GUI components (Zhang et al., 2025b; Cai et al., 2024), or enabling direct post-generation artifact manipulation (Xie et al., 2025; Vaithilingam et al., 2024; Suh et al., 2024; Gero et al., 2022; Petridis et al., 2024; Jiang et al., 2023). Together, these paradigms illustrate diverse strategies for balancing the openness of prompting with the precision of interaction.
@@ 2.2. Human–GenAI Interaction Models and Design Patterns
Interaction design in modern UIs has evolved from early WIMP and direct-manipulation paradigms to richer post-WIMP models that foreground instruments and objects over low-level commands (Van Dam, 1997; Lee et al., 2021). Direct-manipulation interfaces emphasized immediate, visible control—users act on representations of domain objects and observe one-to-one effects (Shneiderman, 1983). Building on this view, Beaudouin-Lafon’s Instrumental Interaction framed instruments as mediators between user actions and artifacts, where each instrumented action maps directly to a target operation (Beaudouin-Lafon, 2000; Beaudouin-Lafon and Mackay, 2000). Subsequent work introduced “substrates” as “places for interaction”, a mechanism to combine power and simplicity by structuring where and how instruments apply in complex interfaces (Mackay and Beaudouin-Lafon, 2025). These classical models are well suited to deterministic manipulation, but they assume a one-to-one relation between an interaction and its effect on an artifact. Generative systems violate that assumption: a single high-level instruction (e.g., a free-form text prompt) can trigger multiple, nonlocal transformations across an artifact or even across multiple artifacts. This multiplicity and indirection require rethinking how instruments, intentions, and objects are modeled in an era of GenAI.
Recent HCI work has begun to explore this space. A series of surveys and taxonomies catalog interaction modalities and recurring patterns. For example, Lehmann et al.  (Lehmann and Buschek, 2024) analyze how UI affordances mediate access to model capabilities, Luera et al.  (Luera et al., 2024) review input modalities across GenAI applications, Gao et al.  (Gao et al., 2024) propose a taxonomy of human–LLM communication modes, and Hu et al.  (Hu et al., 2025) identify general paradigms, i.e., GenAI as medium, tool, partner, and mediator. While valuable as high-level accounts, these taxonomies provide only coarse-grained descriptive and discriminative power and offer limited generative power. For example, Gao et al.  (Gao et al., 2024) label a prompt decomposition subcategory, which is named as “UI for reasoning” in their paper, but cannot distinguish whether decomposition is performed by the human (Wu et al., 2022) or GenAI (Jiang et al., 2023). Following Hu et al.  (Hu et al., 2025), these systems are similarly grouped under “GenAI as medium and UI affordance for exploration”, yet cannot differentiate whether interactions are initiated by the human (Brade et al., 2023) or GenAI (Suh et al., 2024), and whether the human (Suh et al., 2024) or GenAI (Brade et al., 2023) ultimately operates on the artifact. Their main drawback is the lack of fine-grained characterization of interaction flow between humans and AI, leading to the insufficient descriptive and discriminative powers. Shen et al.  (Shen et al., 2025a) advanced them by introducing detailed considerations of action timing and initiators beyond core purposes for interaction, i.e., restricting, expanding, organizing, and refining prompts. However, it still falls short in distinguishing how intent is formulated (e.g., text- vs. artifact-grounded) and what forms the post-interaction information takes (e.g., linguistic vs. non-linguistic). For example, Shen et al.  (Shen et al., 2025a) cannot characterize the difference between Fig. 2-e and Fig. 2-f as “AI-initiated, post-invocation instruction extension”, though they provide different ways to facilitate user intent formulation with pure text prompts or existing artifacts.
Complementing these descriptive efforts, Riche et al.  (Riche et al., 2025) introduce AI-Instruments, extending instrumental interaction (Beaudouin-Lafon, 2000) to generative settings by making prompts themselves manipulable objects and articulating principles of reification, reflection, and grounding. Though these high-level guidelines help consider how human-AI interaction should be designed, they also suffer from limited and coarse-grained descriptive or discriminative power across diverse interfaces. The main reason is their incapability in capturing system-internal information-flow differences, such as interaction timing, the explicit role of artifacts, or how composite instructions combine multiple modalities, as shown in Fig. 2. For example, Riche et al.  (Riche et al., 2025) designed four interface prototypes according to the principles. In their work, “Fillable Brushes” leverage user interactions on existing artifacts and new prompts for next-step AI generation (similar to Fig. 2-d and P4 in our summarized paradigms in Tab. 2). In “Fragments”, AI can break down prompts and generate alternative options based on different design dimensions for users to interact with (similar to P6 in our summarized paradigms in Tab. 2). They both satisfy the three principles, yet clearly correspond to different interaction flow designs that the model cannot structurally distinguish, such as the timing of interactions.
The motivating example (Fig. 2) and the discussion above highlights the need for a unified, systematic model, which can precisely describe how prompts, interactions, and artifacts are combined across various dimensions like timing and resource axes; discriminate between closely related designs; and generate actionable blueprints for evolving real-world interfaces. The next section introduces such a model, addressing these gaps and enabling the descriptive, discriminative, and generative capabilities that prior frameworks lack.